OpenStudy (anonymous):

Simplify the expression (4x − 3)(x + 5).

OpenStudy (anonymous):

4x

OpenStudy (callisto):

I expect you're talking about expanding it (4x − 3)(x + 5) = 4x(x + 5) -3(x + 5) = 4x^2 +5(4x) -3x -3(5) = 4x^2 +20x - 3x -15 = 4x^2 +17x -15